OpenStudy (anonymous):

2. Determine the zeros of f(x) = x3 – 12x2 + 28x – 9.

OpenStudy (anonymous):

Do you know synthetic division?

OpenStudy (anonymous):

By the rational root theorem, possible zeros are +-1, +-3, +-9....

OpenStudy (anonymous):

Yea. Just trying to understand it better

OpenStudy (raden):

9 is one of roots of the polynomial

OpenStudy (anonymous):

this is what i have so far. 1 l 1 -12 28 -9 The remainder is 8 1 -11 17 1 -11 17 8 factors of q= 1, -1 factors of p= 1, -1, 3, -3, 9, -9 Possible rational zeros: 1, -1, 3, -3, 9, -9

OpenStudy (anonymous):

since you were given that 9 is a root, you can find the root of the resulting quadratic after dividing the polynomial by (x-9)... hmm... seems like you only have 1 rational root.

OpenStudy (anonymous):

Thanks